The sum of three numbers in A.P is 39 and their product is 2184. Find the three numbers? Please solve it step by steps for me, as to enable me understand how come about the answer.
Answer by Alan3354(69443)

The sum of three numbers in A.P is 39 and their product is 2184. Find the three numbers? Please solve it step by steps for me, as to enable me understand how come about the answer.
==============
x = middle number
x = 39/3 = 13
(x-a)*x*(x+a) = 2184
(13-a)*13*(13+a) = 2184
(13-a)*(13+a) = 2184/13 = 168
169 - a^2 = 168
a = 1
--> 12, 13 & 14